OpenStudy (anonymous):

Will someone please help! The equation y=-1777x+27,153 can be used to predict the number y of gun deaths in the United States x years after 2000, that is, x=0 corresponds to 2000,x=2 corresponds to 2002, x=5 corresponds to 2005, and so on. Predict the number of gun deaths in 2005 and 2008. In what year will the number of gun deaths be 9383?

OpenStudy (saifoo.khan):

Not again!

OpenStudy (a_clan):

For no. of gun deaths in 2005, x=5 y=-1777*5 + 27153

OpenStudy (anonymous):

>.> 18268?

OpenStudy (saifoo.khan):

y=-1777*8 + 27153 y=12937

OpenStudy (anonymous):

EIGHTEEN THOUSAND GUN DEATHS IN 05 ?!?!?

OpenStudy (saifoo.khan):

alan, u may buy a gun too.

OpenStudy (a_clan):

substitute 9383 in y to get the value of x. add x to 2000, to get the year in which no. of deaths were 9383

OpenStudy (anonymous):

2010

OpenStudy (anonymous):

was last year ^^

OpenStudy (anonymous):

but also is the year 9383 were killed by guns